# Extragalactic Stellar Tidal Streams: Observations meet Simulation

### v1

@@ 7 Conclusions and Outlook
From the results obtained from comparing the stream frequency, characteristics of the DES galaxy sample observations with cosmological simulations, we can conclude that overall the methods applied here work well and provide a valid reference for the analysis of stellar streams.
Generally the predictions of the simulations are in agreement with the results of the analysis carried out on the DES sample, following the same approach to visual inspection as the present work, used as a reference (Miro-Carretero et al., 2024), and with previous work reported in the literature, as presented in Section 6.1. This provides a degree of confidence in the simulation predictions regarding detection of streams in future surveys at surface brightness limits for which we do not have observations today. The cosmological simulations we have have analysed here predict that, in the absence of a confusion limit due to background and foreground sources, and with a pixel size similar to the one of the DECam instrument, a frequency of almost 70% in the detection of streams around galaxies can be achieved for a surface brightness limit of 32 mag $\mathrm{arcsec}^{-2}$. This prediction can be extrapolated to other observations taking into account the effect of the pixel size on the S/N. Nevertheless, there are some noticeable differences in the stream morphologies between observations and simulations and between simulations themselves, that should be further analysed in order to understand their origin and be able to evolve the specific simulation resolution and physics modelling required for stream analysis where needed.
We present a method for comparison of stream observations with cosmological simulations based on novel tools for generation of mock images. When inspecting the mock-images, we follow exactly the same approach and apply the same criteria, with the same team, as we did for inspecting the DES observational image sample, and reported in Miro-Carretero et al. (2024).
In this work, we only compare with the output of the simulations at $z=0$ Exploring the evolution of streams from high redshift to the current time in the simulations could provide more insight into the history, mass ratio and kinematics of the preceding mergers, allowing to compare the simulated reality with the conclusions of the visual inspection. This may also help to understand the differences in stream morphology that have been highlighted in this work.
Overall, the results of our work indicate that surveys reaching a surface brightness limit fainter than 31 mag arcsec-2 would be able to reach a stellar tidal stream detection rate of at least 50%, and thereby test the predictions of the $\Lambda$CDM model as implemented by state-of-the-art cosmological simulations.
@@ Acknowledgements.
JMC wants thank the Leiden Observatory for hosting and providing computer infrastructure and facilities for carrying out part of this work, as well as the Universidad Complutense de Madrid for providing computer infrastructure used in this work. JMC wants to thank Yves Revaz for support in the use of the pNbody tool. DMD acknowledges the grant CNS2022-136017 funding by MICIU/AEI /10.13039/501100011033 and the European Union NextGenerationEU/PRTR and finantial support from the Severo Ochoa Grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 and project (PDI2020-114581GB-C21/ AEI / 10.13039/501100011033). JMC and MAGF acknowledge financial support from the Spanish Ministry of Science and Innovation through the project PID2022-138896NB-C55 APC acknowledges support from a Taiwan Ministry of Education Yushan Fellowship and Taiwan National Science and Technology Council grants 112-2112-M-007-017 and 113-2112-M-007-009. SRF acknowledge support from the Knut and Alice Wallenberg Foundation, the Swedish Research Council (grant 2019-04659), and the Swedish National Space Agency (SNSA Dnr 2023-00164). SRF also acknowledges financial support from the Spanish Ministry of Science and Innovation through the project PID2021-123417ob-i00. For this work we have used GNU Astronomy Utilities (Gnuastro, ascl.net/1801.009) versions $0.17$, $0.18$ and $0.20$. Work on Gnuastro has been funded by the Japanese MEXT scholarship and its Grant-in-Aid for Scientific Research (21244012, 24253003), the European Research Council (ERC) advanced grant 339659-MUSICOS, and from the Spanish Ministry of Economy and Competitiveness (MINECO) under grant number AYA2016-76219-P. M.A acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA and the grant PID2021-124918NA-C43. This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the Ministry of Science and Technology of Taiwan, and National Tsing Hua University. This work is supported by the National Science Center, Poland under Agreement No. 2020/39/B/ST9/03494. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship (grant number MR/V023381/1).

### v2

@@ 8 Conclusions and Outlook
From the results obtained from comparing the stream frequency, characteristics and photometry of the DES galaxy sample observations with cosmological simulations, we can conclude that overall the methods applied here work well and provide a valid reference for the analysis of stellar streams.
Generally the predictions of the simulations are in agreement with the results of the analysis carried out on the DES sample, following the same approach to visual inspection as the present work, used as a reference (Miro-Carretero et al., 2024), and with previous work reported in the literature, as presented in Section 7.1. This provides a degree of confidence in the simulation predictions regarding detection of streams in future surveys at surface brightness limits for which we do not have observations today. The cosmological simulations we have have analysed here predict that, in the absence of a confusion limit due to background and foreground sources, and with a pixel size similar to the one of the DECam instrument, a frequency of almost 70% in the detection of streams around galaxies can be achieved for a surface brightness limit of 32 mag $\mathrm{arcsec}^{-2}$. This prediction can be extrapolated to other observations taking into account the effect of the pixel size on the S/N.
Nevertheless, there are some noticeable differences in the stream morphologies between observations and simulations and between simulations themselves, that should be further analysed in order to understand their origin and be able to evolve the specific simulation resolution and physics modelling required for stream analysis where needed.
We present a method for comparison of stream observations with cosmological simulations based on novel tools for generation of mock images and measurement of photometry parameters. When inspecting the mock-images, we follow exactly the same approach and apply the same criteria, with the same team, as we did for inspecting the DES observational image sample, and reported in Miro-Carretero et al. (2024).
In this work, we only compare with the output of the simulations at $z=0$. Exploring the evolution of streams from high redshift to the current time in the simulations could provide more insight into the history, mass ratio and kinematics of the preceding mergers, allowing to compare the simulated reality with the conclusions of the visual inspection. This may also help to understand the differences in stream morphology that have been highlighted in this work.
Overall, the results of our work indicate that surveys with an instrument of similar characteristics and pixel-size as DECam and, reaching a surface brightness limit fainter than 31 mag arcsec-2 would be required to attain a stellar tidal stream detection rate of at least 50%, and thereby test the predictions of the $\Lambda$CDM model as implemented by state-of-the-art cosmological simulations.
@@ Acknowledgements.
JMC thanks the Leiden Observatory for hosting and providing computer infrastructure and facilities for carrying out part of this work, as well as the Universidad Complutense de Madrid for providing computer infrastructure used in this work. JMC thanks Yves Revaz for support in the use of the pNbody tool. JMC and MAGF acknowledge financial support from the Spanish Ministry of Science and Innovation through the project PID2022-138896NB-C55 DMD acknowledges the grant CNS2022-136017 funding by MICIU/AEI /10.13039/501100011033 and the European Union NextGenerationEU/PRTR and finantial support from the Severo Ochoa Grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 and project (PDI2020-114581GB-C21/ AEI / 10.13039/501100011033). APC acknowledges support from a Taiwan Ministry of Education Yushan Fellowship and Taiwan National Science and Technology Council grants 112-2112-M-007-017 and 113-2112-M-007-009. SRF acknowledge financial support from the Spanish Ministry of Economy and Competitiveness (MINECO) under grant numbers AYA2016-75808-R, AYA2017-90589-REDT, PID2021-123417OB-I00 and S2018/NMT-429. For this work we have used GNU Astronomy Utilities (Gnuastro, ascl.net/1801.009) versions $0.17$, $0.18$ and $0.20$. Work on Gnuastro has been funded by the Japanese MEXT scholarship and its Grant-in-Aid for Scientific Research (21244012, 24253003), the European Research Council (ERC) advanced grant 339659-MUSICOS, and from the Spanish Ministry of Economy and Competitiveness (MINECO) under grant number AYA2016-76219-P. MA acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA and the grant PID2021-124918NA-C43. This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the Ministry of Science and Technology of Taiwan, and National Tsing Hua University. This work is supported by the National Science Center, Poland under Agreement No. 2020/39/B/ST9/03494. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship (grant number MR/V023381/1). CSF acknowledge STFC Consolidated Grant ST/X001075/1 and support from the European Research Council through ERC Advanced Investigator grant, DMIDAS [GA 786910] to CSF. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equip- ment was funded by BEIS capital funding via STFC capi- tal grants ST/K00042X/1, ST/P002293/1 and ST/R002371/1, Durham University and STFC operations grant ST/R000832/1.DiRAC is part of the National e-Infrastructure.
Appendix A Mock-Images Photometry
Figure 19: Histogram of the distribution of average surface brightness measured on the stream for the $g$ (top) $r$ (centre) and $z$ (bottom) bands.
Figure 20: Histogram of the distribution of average $(g-r)_{0}$ colour measured on the stream.
Figure 21: Histogram of the distribution of average distance of the streams to the centre of the host galaxy.